# KAG-Thinker: Interactive Thinking and Deep Reasoning in LLMs via Knowledge-Augmented Generation

### v2

@@ Abstract
In this paper, we introduce KAG-Thinker, which upgrade KAG to a multi-turn interactive thinking and deep reasoning framework powered by a dedicated parameter-light large language model (LLM). Our approach constructs a structured thinking process for solving complex problems, enhancing the the logical coherence and contextual consistency of the reasoning process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. Following the Logical Form guided retrieval and reasoning technology route of KAG, this framework first decomposes complex questions into independently solvable sub-problems (which are also referred to as logical forms) through breadth decomposition. Each such logical form is represented in two equivalent forms—natural language and logical function—and subsequently classified as either a Knowledge Retrieval or Reasoning Analysis task. Dependencies and parameter passing between these tasks are explicitly modeled via logical function interfaces. In the solving process, the Retrieval function performs retrieval tasks. It retrieves one-hop structured and unstructured information of specified knowledge unit. While the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary module to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving module to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn interactive thinking and deep reasoning to align the model with our structured reasoning paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that our model outperforms state-of-the-art models on seven benchmark datasets, achieving an average improvement of 4.1%. In addition, KAG-Thinker not only enhances the reasoning stability of the KAG framework but also achieves performance comparable to the model with 10 times its parameters. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
@@ 1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG) with predefined workflow, the representative works include Lewis et al. (2020)Trivedi et al. (2023), and Li et al. (2025a). These models often use the question to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns to mimic the deeper and more deliberative reasoning as "slow thinking"  Pan et al. (2025), a paradigm inspired by human cognitive processes, to determine the timing of retrieval dynamically. Exemplified by Jin et al. (2025b)Sun et al. (2025), and Wang et al. (2025b). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unreliable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
@@
In order to achieve logical reasoning with more accurate judgments and reduced biases (Li et al., 2025c), recent work has attempted various approaches, such as the atomic reasoning actions decomposition by  Wu et al. (2024) , hierarchical and structured reasoning by  Yang et al. (2025) , and verify reasoning trajectories with symbolic solver by  Guan et al. (2025) . Moreover,  Sumers et al. (2024) propose a cognitive architectures for language agents with three key dimensions: information storage, action space and decision-making procedure. Benefit from these previous works and inspired by Slow Thinking - which solves complex problem with access to additional computational resources, full attention, and sophisticated logical reasoning  Booch et al. (2021) - we propose an innovative structured thinking and reasoning model as shown in Figure 1. It enables deep thinking in LLM by simulating human cognitive mechanisms without reinforcement learning DeepSeek-AI et al. (2025); OpenAI (2024). We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external knowledge retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-question using its own knowledge, we rely on external KBs. However, search results from external KB are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module to extract the core relevant information. To facilitate a clearer understanding, we will introduce these key innovations in detail in the following paragraphs.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. Moreover, predefined reasoning adopts structured and modular RAG pipelines that limit flexibility in evolving and open-ended tasks  Li et al. (2025c, b).
To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025b) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity.  Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.

### v3

@@ Abstract
In this paper, we introduce KAG-Thinker, which upgrade KAG to a multi-turn interactive thinking and deep reasoning framework powered by a dedicated parameter-light large language model (LLM). Our approach constructs a structured thinking process for solving complex problems, enhancing the the logical coherence and contextual consistency of the reasoning process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. Following the Logical Form guided retrieval and reasoning technology route of KAG, this framework first decomposes complex questions into independently solvable sub-problems (which are also referred to as logical forms) through breadth decomposition. Each such logical form is represented in two equivalent forms—natural language and logical function—and subsequently classified as either a Knowledge Retrieval or Reasoning Analysis task. Dependencies and parameter passing between these tasks are explicitly modeled via logical function interfaces. In the solving process, the Retrieval function performs retrieval tasks. It retrieves one-hop structured and unstructured information of specified knowledge unit. While the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary module to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving module to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn interactive thinking and deep reasoning to align the model with our structured reasoning paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that our model outperforms state-of-the-art trained deep search models on seven benchmark datasets, achieving an average improvement of 4.1%. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
@@ 1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG) with predefined workflow, the representative works include Lewis et al. (2020)Trivedi et al. (2023), and Li et al. (2025a). These models often use the question to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns to mimic the deeper and more deliberative reasoning as "slow thinking"  Pan et al. (2025), a paradigm inspired by human cognitive processes, to determine the timing of retrieval dynamically. Exemplified by Jin et al. (2025b)Sun et al. (2025), and Wang et al. (2025c). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unreliable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
@@
In order to achieve logical reasoning with more accurate judgments and reduced biases (Li et al., 2025b), recent work has attempted various approaches, such as the atomic reasoning actions decomposition by  Wu et al. (2024) , hierarchical and structured reasoning by  Yang et al. (2025) , and verify reasoning trajectories with symbolic solver by  Guan et al. (2025) . Moreover,  Sumers et al. (2024) propose a cognitive architectures for language agents with three key dimensions: information storage, action space and decision-making procedure. Benefit from these previous works and inspired by Slow Thinking - which solves complex problem with access to additional computational resources, full attention, and sophisticated logical reasoning  Booch et al. (2021) - we propose an innovative structured thinking and reasoning model as shown in Figure 1. It enables deep thinking in LLM by simulating human cognitive mechanisms without reinforcement learning DeepSeek-AI et al. (2025); OpenAI (2024). We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external knowledge retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-question using its own knowledge, we rely on external KBs. However, search results from external KB are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module to extract the core relevant information. To facilitate a clearer understanding, we will introduce these key innovations in detail in the following paragraphs.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. Moreover, predefined reasoning adopts structured and modular RAG pipelines that limit flexibility in evolving and open-ended tasks  Li et al. (2025b).
To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025c) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity.  Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.